# Is Quadratic Gravity ghost free because the ghost is only virtual?

### v1

@@ 3. Quantization formulas for ghosts
This naive expectation however, may have a problem. It will be shown below, it is not possible to enforce this canonical equal time commutators (3.46) if the coordinates commute
@@ 3.1 A toy quantum mechanical model
It seems that the above prescription can be expressed as follows. The second quantization of the ghost corresponds to skew hermitian operators of the coordinates and impulses. The fundamental postulates of QM states that the operator corresponding to an observable $Q$ should be hermitian. In the present model, the hermiticity is spoiled by an i-i unit in the creation and annihilation operators. For the Pais-Uhlenbeck oscillator, it is $P_{2}$ and $Q_{2}$ which have this unwanted factor. The above procedure corresponds to throwing away the $i$ factor that spoils hermiticity in all calculations. That is, to go from $P_{2}$ and $Q_{2}$ to $iP_{2}$ and $iQ_{2}$. This is in agreement with the postulates of Quantum Mechanics presented in the Solvay Congress in 1927.
@@ 3.2 The path integral formulation
$S=\int d\tau\int d^{3}x\frac{3}{2}\partial_{\mu}\phi\partial_{\mu}\phi+\frac{9\gamma}{2}(\partial_{\mu}\partial_{\mu}\phi)^{2}+V(\phi).$
@@ 4. LSZ reduction for ghost variables
5. The full Stelle action in synchronous
@@
$P^{ij}=-2\sqrt{-g}[G^{ijlm}\dot{K}_{lm}+aK^{ij}K-ag^{ij}K_{lm}K^{lm}+aR^{ij}$
$g^{ij}[(\frac{a}{3}+2b)(~R^{3}-3K_{lm}K^{lm}+K^{2})],$ (5.84)
@@
$G^{ijlm}=a\,\frac{g^{il}g^{jm}+g^{im}g^{jl}}{2}+\left(4b-\frac{a}{3}\right)g^{ij}g^{lm},$
@@
$\Gamma_{pqij}\equiv\frac{1}{a}\frac{g_{pi}g_{qj}+g_{pj}g_{qi}}{2}-\frac{4b-a/3}{12ba}g_{pq}g_{ij}$ (5.85)
@@
$\dot{K}_{ab}=\Gamma_{abij}\bigg\{\frac{P^{ij}}{2\sqrt{-g}}+aK^{ij}K-ag^{ij}K_{lm}K^{lm}+a_{ij}$
$g^{ij}\left[\left(\frac{a}{3}+2b\right)(~R^{3}-3K_{lm}K^{lm}+K^{2}\right]\bigg\},$
@@
[…]
@@
$S_{E}=\int_{\tau_{i}}^{\tau_{f}}d\tau\int d^{3}x\sqrt{g}\left(\frac{a}{2}W_{E}^{2}+bR_{E}^{2}+\frac{2}{\kappa^{2}}R_{E}\right).$ (5.89)
This result is important for the following reason. The theory that is constructed with this quantization is going to be unitary. However, it was not warranted that it will coincide with the euclidean Stelle gravity version. The Stelle theory is known to be renormalizable, and this property would be broken if another lagrangian was to be found. The fact that it coincides means that all the good properties of the Stelle model will be preserved, in particular renormalizability, while keeping unitarity [1].
The extra factor in the last path integral is
@@
6. The Stelle model reconsidered
Consider again the formulas (6.93). These formulas show that the oscillator structure constants corresponding to the free Stelle model are the same as the Pais-Uhlenbeck oscillator found above sections. The task is to understand which new constants does the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ to imaginary values induces.
6.1 About the mapping $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$
The reasoning given above leads to the coefficients $c_{i}$ defining the commutation relations for the creation and annihilation operators in (2.40). However, the quantization prescription requires to make the continuation of the ghost variables to imaginary values, that is, $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. It is natural to ask how (2.40) is changed under this continuation. Assume that it leads to new commutation relations
$A_{p}(k),A_{q}^{c}(k^{\prime})]=d_{1}\delta_{pq}\omega_{1}(k)\delta(k-k^{\prime}),\qquad[B_{p}(k),B_{q}^{c}(k^{\prime})]=d_{2}\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$
$C_{p}(k),C_{q}^{c}(k^{\prime})]=d_{3}\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$ (6.92)
with new constants $d_{i}$ with $i=1,2,3$. The capital letter is to indicate that these creation and annihilation operators are not the same as in (2.40). The task is to understand how to calculate these new constants $d_{i}$. Note that even though $c_{3}=0$, it was not assumed here that $d_{3}=0$, the idea is to be as general as possible in the following discussion.
In the above presentation, one condition was to ensure that
$Q_{1\mu\nu},Q_{2\mu\nu}]=[h_{1\mu\nu},\dot{h}_{2\mu\nu}]=\text{pure gauge terms}.$
This is not the case now, and can be seen as follows. As shown above in (2.36)
$Q_{2\mu\nu},P_{2\mu\nu}]=[h_{1\mu\nu},\dot{h}_{2\mu\nu}]-[P_{1\mu\nu},Q_{1\mu\nu}].$
Since the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ change the sign of the commutator of the left on the last formula, now the identity term of $h_{1\mu\nu},\dot{h}_{2\mu\nu}]-[P_{1\mu\nu},Q_{1\mu\nu}]$ has to have the right sign. This is possible only if now $h_{1\mu\nu},\dot{h}_{2\mu\nu}]$ has a non zero identity term, engineered in such a way that it changes the sign of $P_{1\mu\nu},Q_{1\mu\nu}]$. Therefore now it has to required that both
$P_{1\mu\nu}(x),Q_{1\alpha\beta}(y)]=\frac{i}{2}[(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu})+\text{gauge terms}]\delta(x-y),$
$P_{2\mu\nu}(x),Q_{2\alpha\beta}(y)]=\frac{i}{2}[(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu})+\text{other gauge terms}]\delta(x-y).$
The minus sign of the second commutator is now avoided. This leads to two conditions for the coefficients $d_{i}$.
It is interesting that one condition is still missing in the analysis. There are three coefficients and, up to now, two equations. It may be enforced that $P_{1\mu\nu}(x),P_{2\alpha\beta}(y)]$ is a pure gauge, as before. This was the case in the Pais-Uhlenbeck oscillator, as show in (3.54). Based on this, the same condition may be assumed here. This leads to $d_{3}=0$, again the third mode is not present as an external line. The resulting mapped constants are
$d_{1}=d_{2}=1.$ (6.93)
Again, these are the same as the transformed constants for the Pais-Uhlenbeck oscillator,, described in section 2.
6.2 The LSZ formula for Stelle gravity in detail
By collecting all the information given above, the LSZ rules for the model may be written. They are given by analogy with (4.78)
@@ 6.2 The LSZ formula for Stelle gravity in detail
Here $h^{\mu\nu}_{in}$ is a free field graviton operator of the form described in the previous sections. Despite the similarity with the standard textbook formula, there are differences. The prescription needed to sandwich this operator between the physical states is the crucial point for these differences.
As the quantization prescription requires $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$, it follows that the creation and annihilation operators (6.92) in the field $h_{\mu\nu}^{in}$ in (6.94) have to be employed. In other words, the operators $A_{p}(k)$, $B_{p}(k)$ are the ones inside $h^{in}_{\mu\nu}$, these correspond to the structure constants $d_{i}$ with $i=1,2$. The operator $C_{p}(k)$ is gone, since $d_{3}=0$. The third mode does not participate as an external line. Explicitly the free graviton operator is
$h_{in}^{\mu\nu}=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{1}(k)}[A_{p}(k)\epsilon_{1}^{p\mu\nu}(k)e^{i\omega_{1}(k)t-ik\cdot x}+A^{c}_{p}(k)\epsilon_{1}^{\ast p\mu\nu}(k)e^{-i\omega_{1}(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[B_{p}(k)\epsilon_{2}^{p\mu\nu}(k)e^{i\omega_{2}(k)t-ik\cdot x}+B^{c}_{p}(k)\epsilon_{2}^{\ast p\mu\nu}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
There is no need to do the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ with these operators, they are constructed in such a way that this is automatic. The operators satisfy the algebra
$A_{p}(k),A_{q}^{c}(k^{\prime})]=\delta_{pq}\omega_{1}(k)\delta(k-k^{\prime}),\qquad[B_{p}(k),B_{q}^{c}(k^{\prime})]=\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$
that is, the standard one. The expansion of the exponential in (6.94) leads to terms related to correlation functions of the following form
@@
where $\hat{O}$ is the operator defined in the exponential in (6.94). This operator removes external legs. The full element (6.95) is to be sandwiched between states which a definite momentum. The prescription is that these states have to be constructed in standard fashion, for instance, a state with definite momentum for the free $h_{3in}^{\mu\nu}$ component
$|k_{1},p>_{in}=A^{c}_{p(in)}(k_{3})|0>,\qquad|k_{2},p>_{in}=B^{c}_{p(in)}(k_{3})|0>.$
The tensor product of these states gives multi-particle states.
The use of these states ensures that the $S$ matrix (4.62) is unitary, as they are standard. The part of the Fock space $|k_{3}>$ corresponding the third mode and their tensor produces is decoupled. The unitarity of the $S$ matrix (4.62) is not spoiled. Two modes are interacting and one is decoupled.

### v4

@@ 3.1 The Pais-Uhlenbeck quartic oscillator
This naive expectation however, may have a problem. It will be shown below, it is not possible to enforce this canonical equal time commutators (3.27) if the coordinates commute [42]
@@ 3.3 The path integral formulation
$S=\int d\tau\int d^{3}x[\frac{3}{2}\partial_{\mu}\phi\partial_{\mu}\phi+\frac{9\gamma}{2}(\partial_{\mu}\partial_{\mu}\phi)^{2}+V(\phi)].$
@@ 4. LSZ reduction for ghost variables
5. The full Stelle action in synchronous gauge
@@
$S=-\frac{1}{2\kappa}\int\sqrt{-g}d^{4}x\bigg[R+\frac{\gamma}{2}R^{2}-\frac{\alpha}{2}C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}\bigg].$
@@
$P^{ij}=-2\sqrt{-g}[G^{ijlm}\dot{K}_{lm}+aK^{ij}K-ag^{ij}K_{lm}K^{lm}+\alpha R^{ij}$
$g^{ij}[(\frac{\alpha}{3}+2\gamma)(~R^{3}-3K_{lm}K^{lm}+K^{2})],$ (5.65)
@@
$G^{ijlm}=a\,\frac{g^{il}g^{jm}+g^{im}g^{jl}}{2}+\left(4\gamma-\frac{\alpha}{3}\right)g^{ij}g^{lm},$
@@
$\Gamma_{pqij}\equiv\frac{1}{a}\frac{g_{pi}g_{qj}+g_{pj}g_{qi}}{2}-\frac{4\gamma-\alpha/3}{12\gamma\alpha}g_{pq}g_{ij}$ (5.66)
@@
$\dot{K}_{ab}=\Gamma_{abij}\bigg\{\frac{P^{ij}}{2\sqrt{-g}}+\alpha K^{ij}K-\alpha g^{ij}K_{lm}K^{lm}+a_{ij}$
$g^{ij}\left[\left(\frac{\alpha}{3}+2\gamma\right)(~R^{3}-3K_{lm}K^{lm}+K^{2}\right]\bigg\},$
@@
The integration over $\pi^{ij}$ leads to a Dirac delta that enforces $K_{ij}=-\frac{1}{2}g^{\prime}_{ij}$, where the sign indicates derivative with respect to the euclidean time. The path integral now is related to $D\dot{g}_{ij}$ due to this identification, but a change of variable to $Dg_{ij}$ can be done by multiplying the integration measure by $\prod_{x_{E}}\Delta\tau^{-6}$. The power of $6$ is due to the fact that $\pi^{ij}$, $g_{ij}$, $K_{ij}$ and $P^{ij}$ have a total of $6$ independent components at each spacetime point.
@@
$S_{E}=\int_{\tau_{i}}^{\tau_{f}}d\tau\int d^{3}x\sqrt{g}\left(\frac{\alpha}{2}W_{E}^{2}+\gamma R_{E}^{2}+\frac{2}{\kappa^{2}}R_{E}\right).$ (5.70)
This result is important for the following reason. The theory that is constructed with this quantization is going to be unitary. However, it was not warranted that it will coincide with the euclidean Stelle gravity version. The Stelle theory is known to be renormalizable, and this property would be broken if another lagrangian was to be found. The fact that it coincides means that all the good properties of the Stelle model will be preserved, in particular renormalizability, while keeping unitarity [1]. The extra factor in the last path integral is
@@
6. The LSZ formula for Stelle gravity
6.1 The operator expression of the S matrix
By collecting all the information given above, the LSZ rules for the model may be written. Forgetting for one moment the ghost consider the rules that appear if one consider the standard Stelle theory. The analogy with (4.59) will lead to
@@ 6.1 The operator expression of the S matrix
Note that the operator on the left corresponds to the Stelle lagrangian with a replacement $h^{\alpha\beta}\to\frac{\delta}{\delta J_{\alpha\beta}}$. In order to be consistent with the procedure of taking $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ one must write this lagrangian in terms of the variables $P_{1}$, $P_{2}$, $Q_{1}$ and $Q_{2}$. Then make the continuation. The results of the previous section show that the result is the free euclidean Stelle gravity. The expansion of the exponential in (6.73) leads to terms related to correlation functions of the following form
@@
where $\hat{O}$ is the operator defined in the exponential in (6.73). This operator removes external legs. The full element (6.74) is to be sandwiched between states which a definite momentum, considered as the standard ones.
However, the above result must indicate which are the free fields $h^{in}_{\mu\nu}$. In particular, how they act on the vacuum. In the following, a simpler gauge will be employed to indicate the structure constants of the creation and annihilation operator, assuming that these constants are gauge independent.
6.2 Second quantization of the model
Consider a small perturbation around the Minkowski metric
[…]
Inspired by the methods of [47]-[48], decompose the perturbation as
$\widetilde{h}_{\mu\nu}=\Psi_{\mu\nu}+\frac{\eta_{\mu\nu}}{2}\bigg(\phi+\frac{\Psi}{3}\bigg).$
The component $\Psi_{\mu\nu}$ is of spin two and the field $\phi$ is a scalar. The last expression for the perturbation assumes that the trace of the perturbation is not necessarily zero, and its value is encoded in $\phi$. The Stelle lagrangian, written in terms of the Weyl tensor. may be linealized and it becomes
$L=\frac{1}{2\kappa}\bigg[\frac{1}{4}\partial_{\mu}\Psi^{\alpha\beta}\partial^{\mu}\Psi_{\alpha\beta}-\frac{3}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\partial_{\beta}\Psi^{\alpha\beta}\partial_{\alpha}\phi+\frac{1}{2}\partial_{\alpha}\phi\partial^{\alpha}\Psi\bigg]$
$\frac{\gamma}{2\kappa}\bigg[-\frac{9}{2}\left(\square\phi\right)^{2}+3\partial_{\alpha}\partial_{\beta}\Psi^{\alpha\beta}\square\phi-\frac{3}{2}\square\Psi\square\phi\bigg]+\frac{\alpha}{2\kappa}\left[\frac{1}{4}\square\Psi_{\mu\beta}\square\Psi^{\mu\beta}\right]+L_{int},$ (6.75)
up to interaction terms collected in $L_{int}$ that appear due to the higher order expansion of the curvature and the metric determinant. The equation of motion for $\phi$ is
$3\square\phi-\partial_{\alpha\beta}\Psi^{\alpha\beta}+\frac{1}{2}\square\Psi=\frac{\gamma}{2}\bigg[18\square^{2}\phi-6\square\partial_{\mu}\partial_{\nu}\Psi^{\mu\nu}+3\square^{2}\Psi\bigg],$ (6.76)
while for the spin two perturbation $\Psi_{\mu\nu}$ the equations become
$\square\Psi_{\mu\nu}+2\partial_{\mu}\partial_{\nu}\phi-\eta_{\mu\nu}\square\phi-6\gamma\square\partial_{\mu}\partial_{\nu}\phi+3\gamma\eta_{\mu\nu}\square^{2}\phi-\alpha\square^{2}\Psi_{\mu\nu}=0.$ (6.77)
The choice $\partial_{\mu}\Psi^{\mu\nu}=0$ leads to
$3\square\phi+\frac{1}{2}\square\Psi=\frac{\gamma}{2}\bigg[18\square^{2}\phi+3\square^{2}\Psi\bigg],$
$\square\partial_{\mu}\phi-3\gamma\square^{2}\partial_{\mu}\phi=0.$ (6.78)
This represents $\partial_{\mu}\phi$ as a massless mode plus a massive mode with mass
$m^{2}=\frac{1}{3\gamma}.$
This together with (6.78) means that
$\frac{1}{2}\square\partial_{\mu}\Psi=\frac{3\gamma}{2}\square^{2}\partial_{\mu}\Psi.$ (6.79)
The trace of the equation for $\Psi_{\mu\nu}$ leads to
$\square\Psi+2\square\phi-4\square\phi-6\gamma\square^{2}\phi+12\gamma\square^{2}\phi-\alpha\square^{2}\Psi=0.$
$\square\partial_{\mu}\Psi-\alpha\square^{2}\partial_{\mu}\Psi=0.$ (6.80)
Therefore $\partial_{\mu}\Psi$ contains only a massless mode, as the last two equations (6.79) and (6.80) are non compatible in other case. The equation (6.77) is now
$\square\Psi_{\mu\nu}+2\partial_{\mu}\partial_{\nu}\phi-6\gamma\square\partial_{\mu}\partial_{\nu}\phi-\alpha\square^{2}\Psi_{\mu\nu}=0.$
This, together with (6.78) shows that $\Psi_{\mu\nu}$ is composed by two modes, plus a third one dictated by the massless part of $\phi$. Then
$\square\Psi_{\mu\nu}+2\partial_{\mu}\partial_{\nu}\phi_{m=0}-\alpha\square^{2}\Psi_{\mu\nu}=0.$
Here the index $m=0$ indicates that only the massless mode enters in consideration. One may postulate that $\Psi_{\mu\nu}$ is composed by a mode $\psi_{2\mu\nu}$ with mass $M^{2}=\frac{1}{\alpha}$ plus a further contribution $\Psi_{a\mu\nu}$. This contribution will be such that
$\square\Psi_{a\mu\nu}-\alpha\square^{2}\Psi_{a\mu\nu}=-2\partial_{\mu}\partial_{\nu}\phi_{m=0}.$
As the left is massless, the expression on the right is also massles. Then both vanish, the left by the action of $\square$ and the right consequently. Therefore, the mode for $\phi$ is only massive.
Also (6.80) implies that the trace $\Psi$ is not zero only for the massless mode.
For the scalar mode $\phi$
$\phi(x)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{3}(k)t+ik\cdot x}],$
where the new energy is given by $\omega_{3}(k)=\sqrt{k^{2}+m^{2}}$. For the spin two perturbation the mode expansion becomes
$\Psi_{\mu\nu}(k)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[a_{\mu\nu}(k)e^{i\omega(k)t-ik\cdot x}+a_{\mu\nu}^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[b_{\mu\nu}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{\ast}_{\mu\nu}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
where $\omega(k)=|k|$ and $\omega_{2}(k)=\sqrt{k^{2}+M^{2}}$. Also the gauge and the above found trace conditions lead to
$b_{\mu\nu}(k)\eta^{\mu\nu}=0,\qquad k^{i}a_{\mu i}(k)=\omega a_{\mu 0}(k).\qquad k^{i}b_{\mu i}(k)=\omega_{3}b_{\mu 0}(k).$
The quantum version of these fields are the operators
$\phi(x)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{3}(k)t+ik\cdot x}],$
$\Psi_{\mu\nu}(k)=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[a_{\mu\nu}a_{p}(k)e^{i\omega(k)t-ik\cdot x}+a_{\mu\nu}^{\ast}a_{p}^{c}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[b_{\mu\nu}(k)b_{p}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{\ast}_{\mu\nu}(k)b_{p}^{c}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
The Gauss-Ostrogradsky method, applied to the lagrangian (6.75) leads to the following momentum
$P_{2}=\frac{\gamma}{2\kappa}\bigg[9\square\phi-3\partial_{\alpha}\partial_{\beta}\Psi^{\alpha\beta}+\frac{3}{2}\square\Psi\bigg]$
$P_{1}=\frac{1}{2\kappa}\bigg[3\dot{\phi}+\partial_{\beta}\Psi^{0\beta}+\frac{1}{2}\partial^{0}\Psi\bigg]-\frac{\gamma}{2\kappa}\bigg[9\square\dot{\phi}-3\partial_{\alpha}\partial_{\beta}\dot{\Psi}^{\alpha\beta}+\frac{3}{2}\square\dot{\Psi}\bigg]$
Some terms do not contribute due to the gauge fixing, by employing the Gupta-Bleuler method and taking mean values with the physical states. Taking into account that $Q_{2}=\dot{\phi}$ and that $\phi$ only contains a massive mode, it is impossible that
$P_{2},Q_{2}]=i,\qquad[Q_{1},Q_{2}]=0.$
as the first implies that $\phi,\dot{\phi}]\neq 0$ while the second is $\phi,\dot{\phi}]=0$. It looks like the third mode is gone [^2] This fact confused the author and it was the reason for writing a version that was erroneous.. This alone does not allows to follow the Gauss-Ostrogradsky procedure. However, the error is to apply the Gauss-Ostrogradsky method for this mode alone. The procedure should be applied to the full $h^{\mu\nu}$ momentum.
Therefore, one should keep going on. For $\Psi_{\alpha\beta}$ it follows that
$P_{2\alpha\beta}=-\frac{\alpha}{2\kappa}\frac{1}{2}\square\Psi_{\mu\beta}+\frac{\gamma}{2\kappa}\bigg[3\delta_{0\alpha}\delta_{\beta 0}\square\phi+\frac{3}{2}\eta_{\alpha\beta}\square\phi\bigg]$
$P_{1\alpha\beta}=\frac{1}{2\kappa}\bigg[-\frac{1}{2}\dot{\Psi}_{\alpha\beta}+\delta_{0(\beta}\partial_{\alpha)}\phi-\frac{1}{2}\eta_{\alpha\beta}\dot{\phi}\bigg]+\frac{\alpha}{2\kappa}\frac{1}{2}\square\dot{\Psi}_{\mu\beta}-\frac{\gamma}{2\kappa}\bigg[3\delta_{0\alpha}\delta_{\beta 0}\square\dot{\phi}+\frac{3}{2}\eta_{\alpha\beta}\square\dot{\phi}\bigg]$
Due to the equations of motions given above for $\Psi_{\mu\nu}$, it is seen the massless mode only contributes to $P_{1\alpha\beta}$ while the massive contributes to $P_{2\alpha\beta}$. It is clear that
$Q_{1\alpha\beta},Q_{2\mu\nu}]=[\Psi_{\alpha\beta},\dot{\Psi}_{\mu\nu}]=[\Psi^{1}_{\alpha\beta},\dot{\Psi}^{1}_{\mu\nu}]+[\Psi^{2}_{\alpha\beta},\dot{\Psi}^{2}_{\mu\nu}].$
$P_{1\alpha\beta},Q_{1\mu\nu}]=-\frac{1}{4\kappa}[\dot{\Psi}^{1}_{\alpha\beta},\Psi^{1}_{\mu\nu}],\qquad[P_{2\alpha\beta},Q_{2\mu\nu}]=-\frac{1}{4\kappa}[\Psi^{2}_{\alpha\beta},\dot{\Psi}^{2}_{\mu\nu}]$
$P_{1\alpha\beta},P_{2\mu\nu}]=0.$
All the above conclusions should be interpreted ”up to a gauge terms” and neglecting terms which will vanish when mean values are taken. The conditions for $Q_{1\alpha\beta},Q_{2\mu\nu}]=0$ and that $P_{1\alpha\beta},Q_{1\mu\nu}]$ is the identity up to gauge terms are
$a_{p}(k),a_{q}^{c}(k^{\prime})]=4\kappa\delta_{pq}\omega_{1}(k)\delta(k-k^{\prime}),\qquad[b_{p}(k),b_{q}^{c}(k^{\prime})]=-4\kappa\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$ (6.81)
The full contribution has to take into account $\phi$. However, this field will lead to terms proportional to $\eta_{\alpha\beta}\eta_{\gamma\delta}$ or $k_{\alpha}k_{\beta}k_{\gamma}k_{\delta}$ which are not proportional to the identity $I=\eta_{\alpha(\mu}\eta_{\beta)\nu}$. Therefore, the algebra of creation and annihilation operators of this mode will not make any difference, and it may be assumed that they are standard
$c_{p}(k),c_{q}^{c}(k^{\prime})]=4\kappa\delta_{pq}\omega_{3}(k)\delta(k-k^{\prime}).$ (6.82)
Only $b_{p}$ has a wrong sign. The prescription $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ has the effect analogous to the one explained in the section about the Pais-Uhlenbeck oscillator. It makes the algebra (6.81) standard.